The fact that generally accepted accounting principles allow companies flexibility in choosing between certain allocation methods can make it difficult for a financial analyst to compare periodic performance from firm to firm.

Suppose you were a financial analyst trying to compare the performance of two companies. Company A uses the double-declining-balance depreciation method. Company B uses the straight-line method. You have the following information taken from the 12/31/2021 year-end financial statements for Company B:

Income Statement
Depreciation expense $ 11,500
Balance Sheet
Assets:
Plant and equipment, at cost $ 115,000
Less: Accumulated depreciation (46,000 )
Net $ 69,000


You also determine that all of the assets constituting the plant and equipment of Company B were acquired at the same time, and that all of the $115,000 represents depreciable assets. Also, all of the depreciable assets have the same useful life and residual values are zero.

Required:

1. In order to compare performance with Company A, estimate what B's depreciation expense would have been for 2021 if the double-declining-balance depreciation method had been used by Company B since acquisition of the depreciable assets.
2. If Company B decided to switch depreciation methods in 2021 from the straight line to the double-declining-balance method, prepare the 2021 journal entry to record depreciation for the year, assuming no journal entry for depreciation in 2021 has yet been recorded.

Double-declining balance
Year 1 (2018)
Year 2 (2019)
Year 3 (2020)
Year 4 (2021)

Record the depreciation expense for 2021.

Transaction General Journal Debit Credit
1      
Answer #1

Answer 1.

Asset was acquired on January 2018

Useful life = 10 years

DDb rate = 1/10*2 = 20%

DDB RATE DEPRECIATION EXPENSE NET VALUE OF ASSET
2018 115,000 20% 23,000 92,000
2019 92,000 20% 18,400 73600
2020 73600 20% 14720 58880
2021 58880 20% 11776 47104

B's depreciation expense would have been for 2021 = 11,776

Answer 2.

Net book value at jan 2021 = 115000 - [11500*3] =80500

Remaining useful life = 10 - 3 = 7 years

DDB rate = 1/7*2 = 28.571429%

Depreciation = 80500 * 28.571429% = 23000

JOURNAL ENTRY FOR 2021

DEBIT CREDIT
Depreciation expense $23,000
Accumulated Depreciation    $23,000